The perimeter of rectangle is 30 yards. The width is 12 yards less than twice the length. Find he length and the width of the rectangle.
Answer by pie90010(44)

Let L be the length, and L - 12 to be the width
P = 2L + 2W so
W = L - 12
knowing P = 30 we can say
2L + 2(L-12) = 30
then just combine like terms and end up with
4L - 24 = 30
add 34 to both sides
4L = 54
and divide 4 by each side
L = 13.5
BUT WAIT, that is only the length, they also require the width in the question
so simply plug it in, using the first equation
W = 13.5 - 12
W = 1.5

so the complete answer is
L = 13.5
W = 1.5